OpenStudy (azionne):

12(x - 6) = 2(x - 9) will medal

OpenStudy (anonymous):

Simplifying 12(x + -6) = 2(x + -9) Reorder the terms: 12(-6 + x) = 2(x + -9) (-6 * 12 + x * 12) = 2(x + -9) (-72 + 12x) = 2(x + -9) Reorder the terms: -72 + 12x = 2(-9 + x) -72 + 12x = (-9 * 2 + x * 2) -72 + 12x = (-18 + 2x) Solving -72 + 12x = -18 + 2x Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '-2x' to each side of the equation. -72 + 12x + -2x = -18 + 2x + -2x Combine like terms: 12x + -2x = 10x -72 + 10x = -18 + 2x + -2x Combine like terms: 2x + -2x = 0 -72 + 10x = -18 + 0 -72 + 10x = -18 Add '72' to each side of the equation. -72 + 72 + 10x = -18 + 72 Combine like terms: -72 + 72 = 0 0 + 10x = -18 + 72 10x = -18 + 72 Combine like terms: -18 + 72 = 54 10x = 54 Divide each side by '10'. x = 5.4 Simplifying x = 5.4

OpenStudy (azionne):

answer has to be reduced mixed number

OpenStudy (azionne):

walk me through it

OpenStudy (anonymous):

man i just took 20 min writing that:(

OpenStudy (azionne):

lol this question wont even up for 20

OpenStudy (anonymous):

can you still medal pls?

OpenStudy (anonymous):

just for the effort

OpenStudy (anonymous):

ty